# DIRECT: Deep Active Learning under Imbalance and Label Noise

### 14 Dec 2023 v1

@@ 5 Experiments
(a) Imbalanced CIFAR-10, two classes
(b) Imbalanced CIFAR-10, three classes
(c) Imbalanced CIFAR-100, two classes
(d) Imbalanced CIFAR-100, three classes
(e) Imbalanced SVHN, two classes
(f) Imbalanced SVHN, three classes
Figure 2: Performance of DIRECT vs different baselines across six different dataset settings.
@@ 5.1 Experiment Setup
Our experiment setup follows closely from Zhang et al. (2022). We use the CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009) and SVHN (Netzer et al., 2011) image classification datasets. The original CIFAR-10 and SVHN datasets have 10 classes while the CIFAR-100 dataset has 100 classes. We construct an extremely imbalanced dataset by grouping a large number of classes into one majority class. For example, suppose we have a dataset with $M$ classes. We generate an imbalanced dataset with $K$ classes ($K<M$) by reusing classes $11,...K-$ 1 from the original dataset and combining the rest of the classes $K,...,M$ into a single majority class $K$. Detailed imbalance ratios are shown in Table 1.
Name $K$ $N_{K}$ $\sum_{k=1}^{K-1}N_{k}$ Imb Ratio $\gamma$
CIFAR-10 $2$ $45000$ $5000$$1$.1111.1111.1111
CIFAR-10 $3$ $40000$ $10000$1250.1250.1250.1250
CIFAR-100 $2$ $49500$ $500$0101.0101.0101.0101
CIFAR-100 $3$ $49000$ $1000$0102.0102.0102.0102
SVHN $2$ $68309$ $4948$0724.0724.0724.0724
SVHN $3$ $54448$ $18809$2546.2546.2546.2546
Table 1: Dataset setting for extremely unbalanced scenario. $N_{K}$ denotes the number of images in the majority class while $\sum_{k=1}^{K-1}N_{k}$ is the total number of images in all minority classes. $\epsilon$ is the class imbalance factor defined in Section 3.1.
We compare our algorithm DIRECT against nine baselines: GALAXY (Zhang et al., 2022), SIMILAR (Kothawade et al., 2021), BADGE (Ash et al., 2019), BASE (Emam et al., 2021), BAIT (Ash et al., 2021), Cluster Margin (Citovsky et al., 2021), Confidence Sampling (Settles, 2009), Most Likely Positive (Jiang et al., 2018; Warmuth et al., 2001, 2003) and Random Sampling. The highlight surrounding the baselines represents the standard error plotted as the confidence intervals. Due to computational constraints, we only have single runs for SIMILAR, BADGE, and BAIT. For all other baselines, we average over four runs.
For all experiments, we use PyTorch’s ResNet-18 model pretrained on ImageNet for initialization. We use the Adam optimization algorithm with 500 epochs for each L and the learning rate set to 0.01.
5.2 Results
Synchronous Labeling ($B_{\text{parallel}}=1$). In our first set of experiments, we compare DIRECT against GALAXY in the synchronous labeling setting by setting $B_{\text{parallel}}=1$ in our algorithm. As shown in Figure 2(a), DIRECT outperforms GALAXY by saving an additional $15\%$ of the labeling budget. Compared to random sampling, DIRECT is able to around $90\%$ of annotation cost.
Batch Labeling. We observe that on all six datasets shown in Figure 2, DIRECT performs comparably to GALAXY even when under parallel annotation ($B_{\text{parallel}}=5$). However, we do observe a slight drop in label-efficiency when comparing to synchronous labeling.
6 Future Work
In this paper, we addressed the batch sampling problem of GALAXY by annotating multiple examples in parallel. However, in practice, examples are often annotated asynchronously. We believe this is a valuable future direction and a potential solution is to utilize an asynchronous variant of one-dimensional active learning algorithm.
@@ Acknowledgements
Appendix A.
Lemma .1.
@@
lt="[LOGO]">

### 01 Feb 2024 v2

@@ 5 Experiments
We conduct experiments under two primary setups:
  1. 1.
    Fine-tuning ResNet-18 with vanilla supervised training on imbalanced datasets similar to Zhang et al. (2022).
  2. 2.
    Fine-tuning large pretrained model (CLIP ViT-B32) with semi-supervised training strategies under the LabelBench framework (Zhang et al., 2023a).
In Section 5.1, we document these setups in detail. For both evaluation setups, we first evaluate the performance of DIRECT under the noiseless setting in Section 5.2, showing its superior label-efficiency and ability to accommodate batch labeling. In Section 5.3, we evaluate deep active learning algorithms under a novel setting with both class imbalance and noisy labels. Under this setting, we also include an ablation study of the performance of DIRECT on various levels of label noises. While we highlight many results in this section, see Appendix B for all results.
@@ 5.1 Experiment Setup
ResNet-18 Experiments. ResNet-18 with passive training has been the standard evaluation in existing deep active literature (Ash et al., 2019; Zhang et al., 2022). Our experiment setup utilizes the CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) image classification datasets. The original forms of these datasets are roughly balanced across 9, 10 or 100 classes. We construct an extremely imbalanced dataset by grouping a large number of classes into one majority class. For example, given a balanced dataset above with $M$ classes. We generate an imbalanced dataset with $K$ classes ($K<M$) by the first $K-1$ classes from the original dataset and combining the rest of the classes $K,...,M$ into a single majority class $K$. Imbalance ratios are shown in Table 1.
For neural network training, we utilize the standard passive training on labeled examples with cross entropy loss and Adam optimizer (Kingma & Ba, 2014). The ResNet-18 model (He et al., 2016) is pretrained on ImageNet (Deng et al., 2009) from the PyTorch library. To address data imbalance, we utilize a reweighted cross entropy loss by the inverse frequency of the number of labeled examples in each class. For experiments with label noise, we further add a $10\%$ label smoothing during training (Müller et al., 2019).
LabelBench Experiments. Proposed by Zhang et al. (2023a), LabelBench evaluates active learning performance in a more comprehensive framework. Here, we fine-tune the large pretrained model from CLIP’s ViT-B32 model (Radford et al., 2021). The framework also utilizes semi-supervised learning method FlexMatch (Zhang et al., 2021) to further leverage the unlabeled examples in the pool for training. We conduct experiments on the two imbalanced datasets in LabelBench, with FMoW (Christie et al., 2018) and iWildcam (Beery et al., 2021). Similar to the ResNet-18 experiments, we use a $10\%$ label smoothing in the loss function to improve training under label noise. We did find FlexMatch to perform poorly under the combination of imbalance and label noise, so we used the passive training method for label noise experiments.
Name # Classes $K$ $N$ Imb Ratio $\gamma=\frac{\min_{k}N_{k}}{\max_{k^{\prime}}N_{k^{\prime}}}$
Imb CIFAR-10 $2$ $50000$$1$.1111.1111.1111
Imb CIFAR-10 $3$ $50000$1250.1250.1250.1250
Imb CIFAR-100 $2$ $50000$0101.0101.0101.0101
Imb CIFAR-100 $3$ $50000$0102.0102.0102.0102
Imb CIFAR-100 $10$ $50000$0110.0110.0110.0110
Imb SVHN $2$ $73257$0724.0724.0724.0724
Imb SVHN $3$ $54448$2546.2546.2546.2546
PathMNIST $2$ $89996$1166.1166.1166.1166
FMoW $62$ $76863$0049.0049.0049.0049
iWildCam $14$ $129809$ $4.57\cdot 10^{-5}$
Table 1: Dataset settings for our experiments. $N$ denotes the total number of examples in our dataset. $\gamma$ is the class imbalance ratio defined in Section 3.1.
(a) Imbalanced CIFAR-10, three classes
(b) Imbalanced CIFAR-100, ten classes
(c) Imbalanced SVHN, two classes
(d) PathMNIST, two classes
(e) FMoW, Balanced Pool Accuracy
(f) Imbalanced SVHN, two classes, number of labeled examples in minority class
Figure 3: Performance comparison of DIRECT against other baselines algorithms in the noiseless but imbalanced setting. (a)-(d) are balanced accuracy on ResNet-18 experiments while (e) shows experiment under the LabelBench framework. $B_{\text{parallel}}$ indicates the number of parallel annotations as mentioned in Section 4.2. $B_{\text{parallel}}=1$ is equivalent with sychornous labeling. Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
5.2 Experiments under Imbalance, without Label Noise
For the noiseless experiment on ResNet-18, we compare against nine baselines: GALAXY (Zhang et al., 2022), SIMILAR (Kothawade et al., 2021), BADGE (Ash et al., 2019), BASE (Emam et al., 2021), BAIT (Ash et al., 2021), Cluster Margin (Citovsky et al., 2021), Confidence Sampling (Settles, 2009), Most Likely Positive (Jiang et al., 2018; Warmuth et al., 2001, 2003) and Random Sampling. We briefly distinguish the algorithms into two categories. In particular, both SIMILAR and Most Likely Positive annotate examples that are similar to existing labeled minority examples, thus can significantly annotate a large quantity of minority examples. The rest of the algorithms primarily optimizes for different notions of informativeness such as diversity and uncertainty. For the LabelBench experiments, due to the large dataset and model embedding sizes, we choose a subset of the algorithms that are computationally efficient and among top performers in the ResNet-18 results. In particular, we compared against BADGE, Margin Sampling (Balcan et al., 2006; Kremer et al., 2014), CORESET and GALAXY.
As highlighted in Figures 1(a) and 3(a)-(d), DIRECT consistently and significantly outperforms existing algorithms on the ResNet-18 experiments. In Figures 1(c) and 3(e), we demonstrate the increased label-efficiency is also consistently shown in the LabelBench experiments. Compared to random sampling, DIRECT can save more than 80% of the annotation cost on imbalanced SVHN experiment of Figure 3(c). In terms of class-balancedness, we consistently observe that both Most Likely Positive and SIMILAR annotating greater number of minority class examples, but significantly underperforms in terms of balanced accuracy (an example showin in Figure 3(f)). While Zhang et al. (2022) has already observed this phenomenon, we can further see that DIRECT collects slightly less minority class examples than GALAXY, but outperforms in terms of balanced accuracy. While it is crucial to optimize class-balancedness for better model performance, we see that both extremes of annotating too few and too many minority examples could lead to worse generalization performances. When too few examples are from minority class, the performance of the minority classes could be significantly hindered. When optimized to annotate as many examples from minority class as possible, the algorithm has to tradeoff annotating informative examples to examples it is more certain to be in the minority class. Together, this suggests an intricate balance between the two objectives, generalization performance and class-balancedness.
We would also like to highlight the ability to handle batch labeling. Across our experiments, we see DIRECT outperforms with different amounts of parallel annotation ($B_{\text{parallel}}=1,5$ and $20$), indicating its general effectiveness. This is in comparison to the synchoronous nature of GALAXY, where it is always using $B_{\text{parallel}}=1$. On Figures 1(a) and 4(a), we see that DIRECT outperforms GALAXY with synchronous labeling. Furthermore, in these experiments we also see using $B_{\text{parallel}}=5$ only affects algorithm performances minimally for DIRECT.
(a) Imbalanced CIFAR-10, two classes, 10% label noise
(b) Imbalanced SVHN, two classes, 10% label noise
(c) iWildcam Balanced Pool Accuracy, $10\%$ label noise
Figure 4: Performance of DIRECT against baseline algorithms under $10\%$ label noise. (a)-(b) are balanced accuracy on ResNet-18 experiments while (c) shows results under the LabelBench framework. Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
(a) Imbalanced CIFAR-100, two classes, no label noise
(b) Imbalanced CIFAR-100, two classes, 10% label noise
(c) Imbalanced CIFAR-100, two classes, 15% label noise
Figure 5: Performance of DIRECT against baseline algorithms under different levels of label noise. Results are averaged over four trials and the shaded areas represent standard errors around the average performances.
5.3 Experiments under Imbalance and Label Noise
We conduct novel sets of experiments under both class imbalance and label noise. Here, for both ResNet-18 and LabelBench experiments, we evaluate against all of the algorithms that performed well under the imbalance but noiseless setting above. For all of our experiments, we introduce a fixed percentage of label noise, where the given fraction of the examples’ labels are corrupted to a different class uniformly at random. For most of our experiements with $10\%$ label noise shown in Figure 4, we observe again that DIRECT consistently improves over all baselines including GALAXY. The results are consistent on ResNet-18 and LabelBench setups, and with different $B_{\text{parallel}}$ values, showing DIRECT’s robustness under label noise.
Different Levels of Label Noise As shown in Figures 5(a)-(c) and 1(b), we observe the results on imbalanced CIFAR-100 with two classes across numerous levels of label noise, with $0\%$, $10\%$, $15\%$ and $20\%$ respectively. In fact, the noiseless experiment in Figure 5(a) is the only setting DIRECT slightly underperforms GALAXY in terms of generalization accuracy. However, we see DIRECT becomes more label-efficient under the more practical label noise settings. It is also worth noting that with high label noise of $20\%$, we observe in Figure 1(b) that existing algorithms underperform random sampling. In contrast, DIRECT significantly outperforms random sampling, saving more than $60\%$ of the annotation cost.
6 Conclusion and Future Work
In this paper, we conducted the first study of deep active learning under both class imbalance and label noise. We proposed an algorithm DIRECT that significantly and consistently outperforms existing literature. In this work, we also addressed the batch sampling problem of current SOTA algorithm, GALAXY (Zhang et al., 2022), by annotating multiple examples in parallel. Studying asynchronous labeling could a natural extension of our work. We think a potential solution is to utilize an asynchronous variant of one-dimensional active learning algorithm. In addition, one can further batch the labeling process across different classes to further accommodate an even larger number of parallel annotators.
@@ Acknowledgements
Impact Statement
In the rapidly evolving landscape of machine learning, the efficacy of active learning in addressing data imbalance and label noise is a significant stride towards more robust and equitable AI systems. This research explores how active learning can effectively mitigate the challenges posed by imbalanced datasets and erroneous labels, prevalent in real-world scenarios.
The positive impacts of this research are multifaceted. It enhances the accessibility and utility of machine learning in domains where data imbalance is a common challenge, such as healthcare, finance, and social media analytics. By improving class-balancedness in annotated sets, models trained on these datasets are less biased and more representative of real-world distributions, leading to fairer and more accurate outcomes. Additionally, this research contributes to reducing the time and cost associated with data annotation, which is particularly beneficial in fields where expert annotation is expensive or scarce.
However, if not carefully implemented, active learning strategies could inadvertently introduce new biases or amplify existing ones, particularly in scenarios where the initial data is severely imbalanced or contains deeply ingrained biases. Furthermore, the advanced nature of these techniques may widen the gap between organizations with access to state-of-the-art technology and those without, potentially exacerbating existing inequalities in technology deployment.
Appendix A Equivalent Objective
Lemma A.1.
@@ Appendix A Equivalent Objective
Appendix B All Results
B.1 Noiseless Results under Imbalance
Figure 6: Imbalanced CIFAR-10, two classes.
Figure 7: Imbalanced CIFAR-10, three classes.
Figure 8: Imbalanced CIFAR-100, two classes.
Figure 9: Imbalanced CIFAR-100, three classes.
Figure 10: Imbalanced CIFAR-100, 10 classes.
Figure 11: Imbalanced SVHN, two classes.
Figure 12: Imbalanced SVHN, three classes.
Figure 13: PathMNIST, two classes.
(a) FMoW Balanced Pool Accuracy
(b) iWildcam Balanced Pool Accuracy
Figure 14: LabelBench results in the noiseless setting.
B.2 Label Noise Results under Imbalance
Figure 15: Imbalanced CIFAR-10, two classes, 10% label noise.
Figure 16: Imbalanced CIFAR-10, three classes, 10% label noise.
Figure 17: Imbalanced CIFAR-100, two classes, 10% label noise.
Figure 18: Imbalanced CIFAR-100, two classes, 15% label noise.
Figure 19: Imbalanced CIFAR-100, two classes, 20% label noise.
Figure 20: Imbalanced CIFAR-100, three classes, 10% label noise.
Figure 21: Imbalanced SVHN, two classes, 10% label noise.
Figure 22: Imbalanced SVHN, three classes, 10% label noise.
(a) FMoW Balanced Pool Accuracy
(b) iWildcam Balanced Pool Accuracy
Figure 23: LabelBench results in the 10% label noise setting.